# QMeCha: quantum Monte Carlo package for fermions in embedding environments

### v1

@@ VI.2 Interacting Electron-Positron systems
These two wavefunctions have been used first in ref. 38 to compute the positron affinity and Ps binding energies for a set of atoms and the $e^{+}\cdot(\text{H}_{2}^{2-})$ molecular system. All energies were compared against the most accurate reported values in the literature at the time, either from CI, analytical explicitly correlated methods, or VMC and DMC. As can be seen from the data summarized in Table 1, for the positronium-bound systems: PsB, PsC, PsO, and PsF, the VMC and DMC total energies achieved were lower than any previously reported calculations, establishing new benchmark values, while the PsH and PsLi results are remarkably competitive against larger multi-reference values. On the other hand, there is still room for improvement at VMC level for the weakly bound multireference systems such as $e^{+}\text{Li}$ and $e^{+}\text{Be}$.
Figure 6: Potential energy surfaces of $e^{+}\cdot(\text{H}_{2}^{2-})$, $2\text{H}+\text{Ps}^{-}$ and $\text{H}+\text{H}^{-}+\text{Ps}$ obtained at the DMC level and reported in ref. 38.
Regarding positronic molecules, the same methodology was tested for $e^{+}\cdot(\text{H}_{2}^{2-})$, a particular system characterized by an unique stabilising positron bond interaction between two repelling hydrogen anions 106, 87, 109. The stability of $e^{+}\cdot(\text{H}_{2}^{2-})$ was a matter of debate due to presence of an apparent double-well minima, that at short distances corresponds to a highly delocalized unbound state of $\text{H}_{2}$ and Ps- [109], while at longer distances as displayed in the PES in Fig. 6, shows a local stable positron bounded $e^{+}\cdot(\text{H}_{2}^{2-})$ structure. Compared to the most accurate DMC values [109], the AGP/EPO, and AGP/PMO wavefunction shown only in ref. 38, demonstrated the ability to describe both minima, while previous CI or HF+J VMC approaches could not clearly capture the dissociated $Ps^{-}$ state[109, 38].
@@
Figure 7: a. Positronic (blue) and electronic (red) densities for $2e^{+}\cdot(\text{H}_{3}^{3-})$. b. Electronic (orange) density in $\text{H}_{3}^{+}$. c. Core (red) and valence (orange) electronic densities in $\text{Li}_{3}^{+}$. Data is taken from ref. 39.
@@ VI.3 Embedding approaches for quantum Monte Carlo
The importance of developing such techniques within the framework of QMC methods[247, 248, 40, 42] lies in the possibility of working with any articulated parametric wavefunctions and with multiple environment parametrizations, allowing the exploration of new approaches to include explicit correlation between the environment and the molecular subsystem described through the electronic Schrödinger equation. Furthermore, QMC can be used to obtain accurate reference calculations supporting the development of computationally cheaper approaches.
Figure 8: Interaction energies of the El-QDO water dimer as a function of the distance between the two subsystems. The results are compared to the SAPT2 + 3(CCD) energy components. The vertical dashed line is the equilibrium geometry of the water dimer. The reported F-pEx curve is obtained by removing from the total SAPT energy all the pure exchange contributions[42]. Data taken from ref. 42.
@@
In the El-QDO embedding method[40, 42] the degrees of freedom of the electronic targeted subsystem and those of the drudons in the environment are described through the single comprehensive Hamiltonian in eq. 5, describing explicit many-body correlation effects between the electronic sub-system and the environment. All the degrees of freedom are described through a collective variational ansatz, in eq. 16, that can be easily integrated in the VMC and DMC frameworks.
@@
Figure 9: a. Solvation energy of the water dimer in the embedding cluster of 28 water molecules modelled by QDOs and point charges[59]. b. Variation of the water dimer’s bond energy as a function of the interaction with the cluster of 28 water molecules. In panels c, d, e we display different configurations of the embedding cluster of waters. $R_{min}$ correspond to the minimal distances between the atoms of the dimer and those of the environment. The original geometry for dR = 0.0 Å is taken from ref. 60 (sites.uw.edu/wdbase). Data taken from ref. 42.
For hydrogen-bonded systems, such as the water dimer, we can see in Fig. 8 adapted from ref. 42, that the interaction energy between a QDO-modeled donor water molecule and a full electronic acceptor monomer[42] is a good approximation to a subset of energy components from Symmetry Adapted Perturbation theory SAPT2 + 3(CCD). The reported F-pEx curve is obtained by removing from the total SAPT energy all the pure exchange contributions[42].
@@
Following the study of the interaction energies of molecular and noble gas atoms[40, 42] the El-QDO approach was also used to study the solvation energies of benzene and water dimers, and the variations of their bond energies in cages of water molecules[40, 42]. In Fig. 9 we report the solvation energies and bond energy variations for the water dimer as shown in ref. 42 as a function of the cage distortion from equilibrium. Here the solvation energy is defined, following ref. 262, as the interaction energy of the molecular dimer with the cage $\text{E}_{solv}(\text{D})=\text{E}(\text{D}+\text{cage})-\text{E}(\text{D})-\text{E}(\text{cage})$, where $\text{E}(\text{D})$ and $\text{E}(\text{cage})$ are the energies of the isolated QDO cage and electronic molecular dimer, and $\text{E}(\text{D}+\text{cage})$ is the total energy of the interacting system. The variation of the dimers bond energy, $\Delta\Delta\text{E}=\text{E}_{solv}(\text{D})-\text{E}_{solv}(\text{M}_{1})-\text{E}_{solv}(\text{M}_{2})$, is defined as the difference between the solvation energy of the dimer and those of the two monomers[262].
@@
Both quantities are reported in Figs. 9a and 9b respectively, as a function of the contraction and expansion of the water cage of 28 molecules that surround the electronic water dimer. The contraction and expansion of the cage is obtained using as reference point the geometric centre $\textbf{R}_{C}$ of the nuclei of the electronic subsystem. The vector of the positions of the oxygen atoms $\textbf{R}_{O}$ is modified by a scalar value dR according to the equation $\textbf{R}^{\prime}_{O}=\textbf{R}_{O}(1+\frac{\text{dR}}{\sqrt{|\textbf{R}_{O}|^{2}}})$ along the vector connecting the atoms with $\textbf{R}_{C}$. The Hydrogen pairs of the water molecules are then translated according to shift of their Oxygen atom preserving the intramolecular angles and bonds so that the single molecular structures remain unchanged.
@@
In Figs. 9c, 9d and 9e the structures of the water dimer in the cage of 28 water molecules are displayed for contractions and expansions of dR=[-1.0,0.0,2.0] where the value dR=0.0 Å corresponds to the original MP2 equilibrium geometry from ref. 60. In these panels the values of $\text{R}_{min}$ correspond to the minimum distance between the QDOs in the environment and the atoms of electronic water dimer.
@@
Table 2: Relative runtimes of ortho-Benzyne (in singlet state), Benzene and the Benzene dimer computed in vacum (V) and in cages of four (4W) thirty (30W) or fifty (30W) water molecules. See ref. 40. El calculations are described at the full electronic level while the El-QDO DMC ones are done by substituting the water molecules with the QDO embedding environment. Test calculations have been done with VMC calculations on 56 CPUs using 12 walkers per CPU and 20 bins with 100 steps per block. The time per single block is averaged over 20 blocks. Relative runtime is calculated as the ratio of time per block multiplied by the square of the ratio of the stochastic error $\sigma$, in order to take into account also the differences in the error bars between the cases in vacuum and in the embedding cage. Complete data can be found in ref. 40.
@@
Here, it is crucial to notice that, although the approach treats the drudons of the QDOs as quantum degrees of freedom in the systems (eq. 3), the computational cost of the integration of the embedding environment is negligible with respect to the cost of the integration of the electronic degrees of freedom, as shown in Tab. 2. In this Table we report the computational cost of the DMC calculations for three sets of system, the first that contains one o-benzyne molecule and four water molecule, is described using both full electronic structure and substituting the four water molecules with 4 QDOs (El-QDO). The energy comparisons of the two approaches, reported in ref. 40, show perfect agreement between the two calculations, yet the computational cost of the El-QDO corresponds essentially to that of the single o-benzyne molecule in vacuum (For full comparisons see Fig. 4 of ref. 40).
This is due to the fact that while the wavefunction of the electronic system is usually represented by a determinant and its computational cost is of $\mathcal{O}(N_{e}^{3})$, for the drudonic particles, that are distinguishable due to the asymmetry in the external potential[41], the wavefunction can be written as a simple Slater product whose computational cost is two orders of magnitude smaller.
@@ Acknowledgements.
Matteo Barborini: Conceptualization (lead); Data curation (lead); Formal analysis (lead); Investigation (lead); Methodology (lead); Project Administration (lead); Software (lead); Supervision (lead); Validation (lead); Visualization (lead); Writing – original draft (lead). Jorge Charry: Data curation (equal); Formal analysis (equal); Investigation(equal); Software (equal); Validation (equal); Visualization (equal); Writing – original draft (supporting). Matej Ditte: Data curation (equal); Formal analysis (equal); Investigation(equal); Software (equal); Validation (equal); Visualization (equal); Writing – original draft (supporting). Andronikos Leventis: Software (supporting); Writing – original draft (supporting). Georgios Kafanas: Software (supporting); Writing – original draft (supporting). Alexandre Tkatchenko: Project Administration (supporting); Supervision (supporting); Funding Acquisition (lead); Resources (lead); Writing – original draft (equal).

### v2

@@ VI.2 Interacting Electron-Positron systems
These two wavefunctions have been used first in ref. 47 to compute the positron affinity and Ps binding energies for a set of atoms and the $e^{+}\cdot(\text{H}_{2}^{2-})$ molecular system. All energies were compared against the most accurate reported values in the literature at the time, either from CI, analytical explicitly correlated methods, or VMC and DMC. As can be seen from the data summarized in Table 1, for the positronium-bound systems: PsB, PsC, PsO, and PsF, the VMC and DMC total energies achieved were lower than any previously reported calculations, establishing new benchmark values, while the PsH and PsLi results are remarkably competitive against larger multi-reference values. On the other hand, there is still room for improvement at VMC level for the weakly bound multireference systems such as $e^{+}\text{Li}$ and $e^{+}\text{Be}$.
Figure 7: Potential energy surfaces of $e^{+}\cdot(\text{H}_{2}^{2-})$, $2\text{H}+\text{Ps}^{-}$ and $\text{H}+\text{H}^{-}+\text{Ps}$ obtained at the DMC level and reported in ref. 47. Partially adapted from J. A. Charry Martinez, M. Barborini and A. Tkatchenko J. Chem. Theory Comput. 18, 4, 2267–2280 (2022); licensed under a Creative Commons Attribution ($C$ BY) license.
Regarding positronic molecules, the same methodology was tested for $e^{+}\cdot(\text{H}_{2}^{2-})$, a particular system characterized by an unique stabilising positron bond interaction between two repelling hydrogen anions 117, 98, 120. The stability of $e^{+}\cdot(\text{H}_{2}^{2-})$ was a matter of debate due to presence of an apparent double-well minima, that at short distances corresponds to a highly delocalized unbound state of $\text{H}_{2}$ and Ps- [120], while at longer distances as displayed in the PES in Fig. 7, shows a local stable positron bounded $e^{+}\cdot(\text{H}_{2}^{2-})$ structure. Compared to the most accurate DMC values [120], the AGP/EPO, and AGP/PMO wavefunction shown only in ref. 47, demonstrated the ability to describe both minima, while previous CI or HF+J VMC approaches could not clearly capture the dissociated Ps- state[120, 47].
@@
Figure 8: a. Positronic (blue) and electronic (red) densities for $2e^{+}\cdot(\text{H}_{3}^{3-})$. b. Electronic (orange) density in $\text{H}_{3}^{+}$. c. Core (red) and valence (orange) electronic densities in $\text{Li}_{3}^{+}$. Adapted from J. Charry, F. Moncada, M. Barborini, L. Pedraza-González, M. T. do N. Varella, A. Tkatchenko and A. Reyes, Chem. Sci. 13, 13795-13802 (2022); licensed under a Creative Commons Attribution ($C$ BY NC) license.[48].
@@ VI.3 Embedding approaches for quantum Monte Carlo
The importance of developing such techniques within the framework of QMC methods[258, 259, 49, 51] lies in the possibility of working with any articulated parametric wavefunction and with multiple environments’ parametrizations, allowing the exploration of new approaches to include explicit correlation between the environment and the molecular subsystem described through the electronic Schrödinger equation. Furthermore, QMC can be used to obtain accurate reference calculations supporting the development of computationally cheaper approaches.
Figure 9: Interaction energies of the El-QDO water dimer as a function of the distance between the two subsystems. The results are compared to the SAPT2 + 3(CCD) energy components. The vertical dashed line is the equilibrium geometry of the water dimer. The reported F-pEx curve is obtained by removing from the total SAPT energy all the pure exchange contributions[51]. Data adapted from M. Ditte, M. Barborini and A. Tkatchenko, J. Chem. Theory Comput. 21, 9, 4466–4480 (2025); licensed under a Creative Commons Attribution ($C$ BY) license.[51].
@@
In the El-QDO embedding method[49, 51] the degrees of freedom of the electronic targeted subsystem and those of the drudons in the environment are described through the single comprehensive Hamiltonian in eq. 5, describing explicit many-body correlation effects between the electronic sub-system and the environment, with the exclusion of the short-range interactions that are not captured. All the degrees of freedom are described through a collective variational ansatz, in eq. 17, that can be easily integrated in the VMC and DMC frameworks.
@@
Figure 10: a. Solvation energy of the water dimer in the embedding cluster of 28 water molecules modelled by QDOs and point charges[70]. b. Variation of the water dimer’s bond energy as a function of the interaction with the cluster of 28 water molecules. In panels c, d, e we display different configurations of the embedding cluster of waters. $R_{min}$ correspond to the minimal distances between the atoms of the dimer and those of the environment. The original geometry for dR = 0.0 Å is taken from ref. 71 (sites.uw.edu/wdbase). Data adapted from M. Ditte, M. Barborini and A. Tkatchenko, J. Chem. Theory Comput. 21, 9, 4466–4480 (2025); licensed under a Creative Commons Attribution ($C$ BY) license.[51].
@@
For hydrogen-bonded systems, such as the water dimer, we can see in Fig. 9 adapted from ref. 51, that the interaction energy between a QDO-modeled donor water molecule and a full electronic acceptor monomer[51] is a good approximation to a subset of energy components from Symmetry Adapted Perturbation theory SAPT2 + 3(CCD). The reported F-pEx curve is obtained by removing from the total SAPT energy all the pure exchange contributions[51]. The removal of these exchange contributions comes from the fact that the El-QDO model is still not generalized to reproduce the short-range interaction limit between the environment and the molecular sub-system described at the electronic level. Following the study of the interaction energies of molecular and noble gas atoms[49, 51] the El-QDO approach was also used to study the solvation energies of benzene and water dimers, and the variations of their bond energies in cages of water molecules[49, 51]. In Fig. 10 we report the solvation energies and bond energy variations for the water dimer as shown in ref. 51 as a function of the cage distortion from equilibrium. Here the solvation energy is defined, following ref. 273, as the interaction energy of the molecular dimer with the cage $\text{E}_{solv}(\text{D})=\text{E}(\text{D}+\text{cage})-\text{E}(\text{D})-\text{E}(\text{cage})$, where $\text{E}(\text{D})$ and $\text{E}(\text{cage})$ are the energies of the isolated QDO cage and electronic molecular dimer, and $\text{E}(\text{D}+\text{cage})$ is the total energy of the interacting system. The variation of the dimers bond energy, $\Delta\Delta\text{E}=\text{E}_{solv}(\text{D})-\text{E}_{solv}(\text{M}_{1})-\text{E}_{solv}(\text{M}_{2})$, is defined as the difference between the solvation energy of the dimer and those of the two monomers[273].
@@
Both quantities are reported in Figs. 10a and 10b respectively, as a function of the contraction and expansion of the water cage of 28 molecules that surround the electronic water dimer. The contraction and expansion of the cage is obtained using as reference point the geometric centre $\textbf{R}_{C}$ of the nuclei of the electronic subsystem. The vector of the positions of the oxygen atoms $\textbf{R}_{O}$ is modified by a scalar value dR according to the equation $\textbf{R}^{\prime}_{O}=\textbf{R}_{O}(1+\frac{\text{dR}}{\sqrt{|\textbf{R}_{O}|^{2}}})$ along the vector connecting the atoms with $\textbf{R}_{C}$. The Hydrogen pairs of the water molecules are then translated according to shift of their Oxygen atom preserving the intramolecular angles and bonds so that the single molecular structures remain unchanged.
@@
In Figs. 10c, 10d and 10e the structures of the water dimer in the cage of 28 water molecules are displayed for contractions and expansions of dR=[-1.0,0.0,2.0] where the value dR=0.0 Å corresponds to the original MP2 equilibrium geometry from ref. 71. In these panels the values of $\text{R}_{min}$ correspond to the minimum distance between the QDOs in the environment and the atoms of electronic water dimer.
@@
Table 2: Relative runtimes of ortho-Benzyne (in singlet state), Benzene and the Benzene dimer computed in vacum (V) and in cages of four (4W) thirty (30W) or fifty (50W) water molecules. See ref. 49. El calculations are described at the full electronic level while the El-QDO DMC ones are done by substituting the water molecules with the QDO embedding environment. Test calculations have been done with VMC calculations on 56 CPUs using 12 walkers per CPU and 20 bins with 100 steps per block. The time per single block is averaged over 20 blocks. Relative runtime is calculated as the ratio of time per block multiplied by the square of the ratio of the stochastic error $\sigma$, in order to take into account also the differences in the error bars between the cases in vacuum and in the embedding cage. Data adapted from M. Ditte, M. Barborini, L. Medrano Sandonas, and A. Tkatchenko Phys. Rev. Lett. 131, 228001 (2023); licensed under a Creative Commons Attribution ($C$) license.[49]
@@
Here, it is crucial to notice that, although the approach treats the drudons of the QDOs as quantum degrees of freedom in the systems (eq. 3), the computational cost of the integration of the embedding environment is negligible with respect to the cost of the integration of the electronic degrees of freedom, as shown in Tab. 2. In this table we report the computational cost of the DMC calculations for three sets of system, the first that contains one o-benzyne molecule and four water molecule, is described using both full electronic structure and substituting the four water molecules with 4 QDOs (El-QDO). The energy comparisons of the two approaches, reported in ref. 49, show perfect agreement between the two calculations, yet the computational cost of the El-QDO corresponds essentially to that of the single o-benzyne molecule in vacuum (For full comparisons see Fig. 4 of ref. 49).
This is due to the fact that while the wavefunction of the electronic system is usually represented by a determinant and its computational cost is of $\mathcal{O}(N_{e}^{3})$, for the drudonic particles, that are distinguishable due to the asymmetry in the external potential[50], the wavefunction can be written as a simple Hartree product whose computational cost is two orders of magnitude smaller.
@@ Acknowledgements.
Matteo Barborini: Conceptualization (lead); Data curation (lead); Formal analysis (lead); Investigation (lead); Methodology (lead); Project Administration (lead); Software (lead); Supervision (lead); Validation (lead); Visualization (lead); Writing – original draft (lead). Jorge Charry: Data curation (equal); Formal analysis (equal); Investigation(equal); Software (equal); Validation (equal); Visualization (equal); Writing – original draft (supporting). Matej Ditte: Data curation (equal); Formal analysis (equal); Investigation(equal); Software (equal); Validation (equal); Visualization (equal); Writing – original draft (supporting). Andronikos Leventis: Software (supporting); Writing – original draft (supporting). Georgios Kafanas: Software (supporting); Writing – original draft (supporting). Alexandre Tkatchenko: Project Administration (supporting); Supervision (supporting); Funding Acquisition (lead); Resources (lead); Writing – original draft (supporting).